OpenStudy (anonymous):

Factor 8x^2-10xy-3y^2

OpenStudy (anonymous):

So what is the curriculum? What course is this?

OpenStudy (anonymous):

Is there a particular method that you must use? Are you not allowed to use the quadratic formula?

OpenStudy (anonymous):

what's the problem?

OpenStudy (anonymous):

factor the trinomial 8x^2-10xy-3y^2

OpenStudy (anonymous):

Consider this: \[ax^2+bx+c\]Where \[a=8,b=-10y, c=-3y^2\]

OpenStudy (anonymous):

Ok what would be the next steps

OpenStudy (anonymous):

multiply 8 and 3?

OpenStudy (anonymous):

\[\Large x = \frac{-b\pm \sqrt{ b^2-4ac }}{2a}\]

OpenStudy (anonymous):

That is the quadratic formula. You could also try splitting the middle.

OpenStudy (anonymous):

that is what i am learning quadratic

OpenStudy (anonymous):

Okay, so solve for \(x\) in terms of \(y\) using the quadratic formula. Do that first.

OpenStudy (anonymous):

my answer is (2x+y)(4x-3)

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Nope, it's not correct.

OpenStudy (anonymous):

(2x-3)(4x+y) ??

OpenStudy (anonymous):

Almost.

OpenStudy (anonymous):

where am I messing up at?

OpenStudy (anonymous):

Simplifying 8x2 + -10xy + -3y2 Reorder the terms: -10xy + 8x2 + -3y2 Factor a trinomial. (2x + -3y)(4x + y) Final result: (2x + -3y)(4x + y)

OpenStudy (anonymous):

You're missing a y for the first factor.

OpenStudy (anonymous):

I dont get where the Y would come from. it was not a GCF

OpenStudy (anonymous):

8x^2 = (4x)(2x) -3y^2 = (-3y)(y)

OpenStudy (anonymous):

8x^2-10xy-3y^2 = 8x^2 - 12xy + 2xy -3y^2 = 8x^2 + 2xy - 12xy -3y^2 = 2x ( 4x + y) - 3y (4x+y) = ...

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

You're welcome. Hope that you understand where the y comes from.